# KVmix: Gradient-Based Layer Importance-Aware Mixed-Precision Quantization for KV Cache

### v1

@@ Abstract
The high memory demands of the Key-Value (KV) Cache during the inference of Large Language Models (LLMs) severely restrict their deployment in resource-constrained platforms. Quantization can effectively alleviate the memory pressure caused by KV Cache. However, existing methods either rely on static one-size-fits-all precision allocation or fail to dynamically prioritize critical KV in long-context tasks, forcing memory-accuracy-throughput tradeoffs. In this work, we propose a novel mixed-precision quantization method for KV Cache named KVmix. KVmix leverages gradient-based importance analysis to evaluate how individual Key and Value projection matrices affect the model loss, enabling layer-specific bit-width allocation for mix-precision quantization. It dynamically prioritizes higher precision for important layers while aggressively quantizing less influential ones, achieving a tunable balance between accuracy and efficiency. KVmix also introduces a dynamic long-context optimization strategy that adaptively keeps full-precision KV pairs for recent pivotal tokens and compresses older ones, achieving high-quality sequence generation with low memory usage. Additionally, KVmix provides efficient low-bit quantization and CUDA kernels to optimize computational overhead. On LLMs such as Llama and Mistral, KVmix achieves near-lossless inference performance with extremely low quantization configuration (Key 2.19bit Value 2.38bit), while delivering a remarkable 4.9× memory compression and a 5.3× speedup in inference throughput.
1 Introduction
@@
Large Language Models (LLMs) [1], such as GPT [2], Llama [3], and their derivatives, have significantly advanced the field of Natural Language Processing (NLP). These models exhibit outstanding performance [4, 5] across a diverse array of tasks, including text generation, question answering, and machine translation. The Key-Value (KV) Cache plays an essential role in the autoregressive decoding process of LLMs. The KV Cache substantially reduces redundant computations in the attention mechanism by storing Key and Value states from preceding time steps for subsequent token generation, thereby significantly accelerating inference efficiency [6]. Nevertheless, as sequence lengths grow, the memory footprint of the KV Cache increases linearly, presenting a formidable challenge to hardware resources. For example, a model with 70 billion parameters may demand over 50GB of memory to maintain the KV Cache when generating a sequence of 20k tokens—a requirement that far exceeds the memory capacity of typical GPUs. In scenarios involving multiple concurrent requests, the KV Cache for each request cannot be shared due to its dependence on unique preceding prompts. Although model parameters can be reused, memory quickly becomes saturated due to the extensive KV Cache demands. Once memory is depleted, data is offloaded to system memory or even disks, resulting in frequent High Bandwidth Memory (HBM) exchanges with system memory. This process causes latency to surge exponentially, leading to catastrophic performance degradation.
@@
The characteristics of the KV Cache outlined above severely restrict the practical deployment and inference efficiency of LLMs in resource-constrained environments, highlighting the pressing need for efficient memory optimization strategies [7, 8]. In recent years, research tackling this issue has predominantly focused on reducing the memory overhead of the KV Cache through quantization and sparsification techniques [9, 10, 11]. Quantization methods, in particular, have gained widespread adoption in industry, significantly contributing to the scalability and accessibility of large-scale models [12]. By quantizing the KV Cache, memory usage can be markedly reduced—for instance, transitioning from FP16 to INT8 halves memory requirements while accelerating memory access and improving throughput. Existing quantization methods have demonstrated impressive model performance even at very low bit-widths. However, these methods either rely on static one-size-fits-all precision allocation schemes [13, 14], lacking flexibility and performance-aware adaptation capabilities, or suffer from the high computational overhead of dynamic quantization while failing to adaptively prioritize critical KV pairs in long-context tasks [15, 16]. Therefore, they are forced to make suboptimal trade-offs among memory usage, model accuracy, and computational throughput.
To address these problems, this paper proposes KVmix, a novel mixed-precision quantization method for KV Cache. KVmix evaluates KV importance using gradient norms of Key and Value projection weights relative to the model’s loss function, enabling layer-specific mixed-precision quantization with flexible bit-width allocation. It proposes a dynamic long-context optimization strategy, keeping full-precision KV pairs for recent pivotal tokens while compressing older pairs. Additionally, KVmix offers low-bit compression strategies and efficient CUDA implementations. KVmix maximizes KV Cache compression and model throughput while maintaining controlled precision. Our specific contributions are as follows:
@@
  • We design efficient CUDA implementations and a high-compression 3-bit quantization method for KVmix. Extensive experimental results show that KVmix achieves nearly lossless model accuracy across multiple LLMs and datasets, as well as a 4.9$\times$ reduction in runtime memory usage and a 5.3$\times$ speedup in inference efficiency, outperforming prior state-of-the-art (SOTA) quantization methods for KV Cache.
2 Related Work and Motivation
2.1 Related Work
@@
To mitigate the memory challenges associated with KV Cache, researchers have developed a variety of optimization approaches, primarily centered on compression techniques and dynamic memory management [17, 18]. We mainly discuss the KV Cache compression techniques related to this work. Existing compression approaches encompass quantization, sparsification, and KV Cache sharing [19, 20, 21].
@@
Extensive research has focused on reducing the memory overhead of KV Cache through quantization. For example, KIVI [13] introduced a 2-bit asymmetric quantization technique, employing per-channel quantization for Keys and per-token quantization for Values. This method substantially improves model throughput while preserving accuracy at low bit-widths. KVQuant [22] proposed a non-uniform quantization strategy, integrating pre-RoPE per-channel Key quantization with per-token Value quantization. It employs offline calibration to manage outliers, achieving robust performance in long-context inference scenarios. QAQ [15] developed a dynamic mixed-precision quantization method that calculates quantization bits for Keys and Values online and optimizes the trade-off between accuracy and compression ratio by predicting attention scores. Atom [23] investigated a mixed-precision scheme involving 4-bit and 8-bit activations, dynamically quantizing activations to adapt to input distributions. SKVQ [16] implemented a sliding window quantization method, maintaining high precision for recent KV pairs while reducing quantization errors through channel reordering. IntactKV [14] simplified the quantization process by preserving critical tokens (e.g., initial tokens) with full precision, thereby sustaining generation quality. WKVQuant [24] further compressed the KV Cache by incorporating weight quantization. Numerous other studies have also made significant contributions to KV Cache quantization [25, 26, 27]. These efforts generally aim to quantize as many KVs as possible to the lowest feasible bit-width. While methods such as KVQuant, QAQ, and Atom have introduced mixed-precision quantization for KV, their primary focus is on addressing outliers that are sensitive to precision changes. In contrast, this work proposes a more flexible quantization framework, enabling users to balance average quantization bit-width and model accuracy during deployment based on their specific needs.
In addition to quantization, sparsification techniques have been widely used to optimize KV Cache. These approaches typically leverage importance assessments derived from the attention weight to dynamically adjust KV storage requirements. H2O [28] introduced a resampling-based strategy that retains KV entries with high attention scores for efficient memory compression. Scissorhands [29] capitalized on the "importance persistence" hypothesis to dynamically evict low-importance KVs, which further improves sparsification efficiency. SnapKV [30] employed predictive modeling of future attention patterns to preemptively discard irrelevant KV entries, while PyramidInfer [31] extended H2O by recognizing that the number of critical KV entries diminishes with increasing model depth, progressively increasing sparsification across layers. We observe that evicting KV from shallow layers propagates the effect to deeper layers through hidden states. Unlike PyramidInfer, we use the contribution of KV to the model output to evaluate the importance of different layers. Our method is orthogonal to existing sparsification methods and can also be used as a guideline for KV sparsification to achieve more accurate KV eviction.
@@
Figure 1: Projection matrix weights of K and V across different layers for the Llama 2-7B model. The heatmap visualizes the projection weights by sampling every sixth element. Additionally, the weight values along the axes have been normalized. "Norm" represents the L2 norm of the weight matrix for each layer, while "Range" indicates the range of values within each layer’s weight matrix.
2.2 Motivation
Current KV Cache quantization methods depend on fixed quantization strategies, which uniformly apply the same bit-width to the Keys or Values across all layers of a model. These methods fail to account for the differing contributions of Keys and Values from various layers to the final output. For each layer of the KV Cache, the computation process is outlined as follows: at time step $t$ and for the $i$-th layer, the input comprises hidden states (denoted as $H_{i-1,t}$), which are supplied by the output of the preceding layer. These hidden states are then utilized to compute the K and V for the current token: $K_{i,t}=W_{k_{i}}\cdot H_{i-1,t}$ and $V_{i,t}=W_{v_{i}}\cdot H_{i-1,t}$, where $W_{k_{i}}$ and $W_{v_{i}}$ represent the projection weights for the K and V, respectively, at the $i$-th layer. Subsequently, the computed $K_{i,t}$ and $V_{i,t}$ are concatenated with the previously stored KV Cache, yielding the complete K and V sequences up to time step $t$: $K_{i,1:t}=[K_{i,1},K_{i,2},\dots,K_{i,t}]$ and $V_{i,1:t}=[V_{i,1},V_{i,2},\dots,V_{i,t}]$. This computation process reveals that the projection weights $W_{k_{i}}$ and $W_{v_{i}}$ determine how the model extracts K and V information from the hidden states. These weights directly influence the quality of the KV generated within the attention mechanism, thereby forming the foundation of the layer’s computation. As a result, the weights of the KV affect the layer’s contribution to the model’s overall output.
To substantiate this analysis, Fig. 1 provides heatmaps of the projection weights $W_{k}$ and $W_{v}$ for the Llama 2-7B model [32], with additional heatmaps for other models included in Appendix A. Examination of these heatmaps yields two critical insights: ① There is considerable variation in the weight values of KV across different layers of the model.Within the same layer, the Key and Value weights display distinct distributional patterns. These observations highlight the fact that the influence of Keys and Values from different layers on the model’s final output is not uniform.
3 Methodology
3.1 KV Importance Analysis
@@
In the self-attention mechanism, the Key and Value are computed by applying linear transformations to the hidden states of the previous layer using KV projection weight matrices and are subsequently combined with the Query to compute the attention output. Consequently, the magnitude of KV projection weights alone is insufficient to measure the importance of the Key and Value of each layer, and a more precise evaluation metric is needed. To address this problem, we propose the KVmix profiler, a gradient-based method that quantifies the contribution of each layer’s Key and Value to the model’s output, enabling a judicious mixed-precision quantization strategy. Specifically, we compute the $L2$ norm of the gradients of the model’s loss function with respect to the Key and Value projection weight matrices for each model layer, and evaluate the importance of the Key and Value components based on these $L2$ norm values. KVmix profiler captures the dynamic sensitivity of these parameters during the model inference process, and provides a layer-specific importance metric to support efficient mixed-precision quantization in subsequent inference stages.
@@
The implementation of KVmix profiler consists of the following three key steps: ① Data preparation and forward propagation. A full-precision model is loaded, and multiple prompts are randomly sampled from a target dataset to serve as input data. These prompts are tokenized into input tensors using the tokenizer. Leveraging the autoregressive property of LLMs, each input tensor is shifted left by one position to be used as the corresponding label tensor for computing the model’s loss function. Subsequently, the loss value for each input is determined through forward propagation. ② Gradient calculation and importance assessment. For the $i$-layer of the model, the gradients of the losses with respect to the projection weights of the Key ($W_{k_{i}}$) and Value ($W_{v_{i}}$) are computed independently. This process begins with backpropagation to calculate the gradients, i.e., $\nabla_{W_{k_{i}}}\mathcal{L}$ and $\nabla_{W_{v_{i}}}\mathcal{L}$, where $\mathcal{L}$ represents the loss function. The magnitude of these gradients is then evaluated using the $L2$ norm, i.e., $\|\nabla_{W_{k_{i}}}\mathcal{L}\|_{2}$ and $\|\nabla_{W_{v_{i}}}\mathcal{L}\|_{2}$. The importance scores of the Keys and Values for each layer can be expressed as:
$s_{k_{i}}=\|\nabla_{W_{k_{i}}}\mathcal{L}\|_{2},s_{v_{i}}=\|\nabla_{W_{v_{i}}}\mathcal{L}\|_{2}$ (1)
@@
A larger $s_{k_{i}}$ or $s_{v_{i}}$ signifies a greater impact of that $i$-th layer’s Key or Value on model’s output. To enhance assessment reliability, the gradient norms can be averaged across multiple prompts ($p$), yielding an average importance score for each layer’s Key and Value:
@@
Using the importance scores, we classify the importance of the Key and Value components across all model layers. The top 20% of the layers of $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ are quantized with high-bit representations (e.g., 3-bit or 4-bit), while the remaining 80% of layers adopt more aggressive low-bit quantization (e.g., 2-bit). This 20%-80% split is not fixed and can be dynamically adjusted according to the requirements to balance the trade-off between model accuracy and memory usage. Increasing the proportion of low-bit quantization layers can further reduce the memory footprint of the KV Cache, but may sacrifice some accuracy. ③Model Configuration and Inference. The KV quantization configuration results derived from the above steps are incorporated into the model configuration, enabling the quantized model to be used for inference. The detailed workflow is depicted in Fig. 2, with the algorithmic procedure outlined in Algorithm B of Appendix B. The profiling is performed offline and therefore does not affect the inference efficiency. In addition, it is once-and-for-all for the inference of the model.
@@
3.2 Asymmetric Low-Bit Quantization
Asymmetric Quantization Strategy. We use per-channel and per-token grouping quantization methods for Key and Value, respectively. The KV Cache has the shape $[B,nh,T,D]$, where $B$ is the batch size, $nh$ is the number of attention heads, $T$ is the token sequence length, and $D$ is the head dimension. When the Key is quantized per channel ($D$), the tensor is reshaped to $[B\times nh\times D,T]$, with each group comprising all tokens of a single channel. This approach is inspired by the distributional properties of the Key Cache that exhibit significant outliers in the channel dimension, i.e., certain channels exhibit significantly large magnitude values. Per-channel quantization isolates errors within each channel and prevents outliers from affecting other channels. When Value is quantized per token, the shape of the tensor is preserved, and each group contains all channels of a single token. Unlike the Key Cache, the Value Cache has no pronounced outliers, but plays a critical role in computing the attention output. Per-token quantization confines errors to individual tokens, preserving the integrity of other important tokens. This asymmetric quantization strategy effectively reduces errors introduced during KV Cache quantization.
Group-Wise Low-Bit Quantization. To minimize the memory usage of the quantized KV Cache, we apply group-wise lowest possible bit quantization for KV. The quantization process is detailed as follows: ① Calculation of scaling factor. The input elements are divided into groups based on the asymmetric quantization strategy (e.g., per-channel for Key or per-token for Value). For each group, we first determine the minimum ($min\_val$) and maximum ($max\_val$) values. The scaling factor $s$ is then calculated as: $s=\frac{max\_val-min\_val}{q_{max}}$, where $q_{max}$ denotes the maximum quantized value. ② Element quantization. Each element is quantized using: $q=round\left(\frac{x-min\_val}{s}\right)$, where $x$ represents the original element value, and $q$ is the quantized value. ③ Clipping of quantized values. The quantized values are constrained to the range $[0,q_{max}]$ through clipping: $q=\max(0,\min(q,{q_{max}}))$. ④ Storage and dequantization. The quantized values are stored using bit operations within a 32-bit integer ($int32$). For 4-bit, 2-bit, and 1-bit quantization, the number of elements in per $int32$ is: $feat\_per\_int=32/bit$. During inference, dequantization is performed using: $x=q\cdot s+min\_val$. For 3-bit quantization, we introduce a new packing strategy to maximize memory efficiency. We organize the quantized elements into blocks of 11, each stored in a 32-bit integer, with the first 10 elements quantized to 3 bits and the 11th element to 2 bits. The clipping range is adjusted based on the element index:
@@ 3.2 Asymmetric Low-Bit Quantization
where $i$ is the element index within a block. This method provides a 10% increase in packing density compared to a uniform 3-bit quantization that can only hold 10 elements per $int32$.
3.3 Dynamic Pivotal Context Selection

### v2

@@ Abstract
The high memory demands of the Key-Value (KV) Cache during the inference of Large Language Models (LLMs) severely restrict their deployment in resource-constrained platforms. Quantization can effectively alleviate the memory pressure caused by KV Cache. However, existing methods either rely on static one-size-fits-all precision allocation or fail to dynamically prioritize critical KV in long-context tasks, forcing memory-accuracy-throughput tradeoffs. In this work, we propose a novel mixed-precision quantization method for KV Cache named KVmix. KVmix leverages gradient-based importance analysis to evaluate how individual Key and Value projection matrices affect the model loss, enabling layer-specific bit-width allocation for mix-precision quantization. It dynamically prioritizes higher precision for important layers while aggressively quantizing less influential ones, achieving a tunable balance between accuracy and efficiency. KVmix introduces a dynamic long-context optimization strategy that adaptively keeps full-precision KV pairs for recent pivotal tokens and compresses older ones, achieving high-quality sequence generation with low memory usage. Additionally, KVmix provides efficient low-bit quantization and CUDA kernels to optimize computational overhead. On LLMs such as Llama and Mistral, KVmix achieves near-lossless inference performance with extremely low quantization configuration (Key 2.19bit Value 2.38bit), while delivering a remarkable 4.9× memory compression and a 5.3× speedup in inference throughput.
Codehttps://github.com/LfLab-AI/KVmix
Introduction
@@
Large Language Models (LLMs) (Vaswani et al. 2017), such as GPT (Radford et al. 2019), Llama (Touvron et al. 2023a), and their derivatives, have significantly advanced the field of Natural Language Processing (NLP). These models exhibit outstanding performance (Hadi et al. 2023; Chang et al. 2024) across a diverse array of tasks, including text generation, question answering, and machine translation. The Key-Value (KV) Cache plays an essential role in the autoregressive decoding process of LLMs. The KV Cache substantially reduces redundant computations in the attention mechanism by storing KV states from preceding time steps for subsequent token generation (Xiao et al. 2023). Nevertheless, as sequence lengths grow, the memory footprint of the KV Cache increases linearly, presenting a formidable challenge to hardware resources. For instance, a 70B-parameters model may require over 50GB of memory to maintain the KV Cache for a 20k-token sequence, exceeding typical GPU capacity. In scenarios involving multiple concurrent requests, the KV Cache for each request cannot be shared due to its dependence on unique preceding prompts. Although model parameters can be reused, memory quickly becomes saturated due to the KV Cache demands. Once memory is depleted, data is offloaded to system memory or even disks, resulting in frequent High Bandwidth Memory (HBM) exchanges with system memory. This process causes latency to surge exponentially, leading to catastrophic performance degradation.
@@
The KV Cache’s characteristics outlined above severely restrict LLM deployment and inference efficiency in resource-constrained environments, underscoring the pressing need for efficient memory optimization (Liu et al. 2024b). Recent research tackling this issue has predominantly focused on reducing the memory overhead of the KV Cache through quantization and sparsification techniques (Shi et al. 2024; Adnan et al. 2024). Quantization methods, in particular, have gained widespread adoption in industry, significantly contributing to the scalability and accessibility of large-scale models (Kumar 2024). Quantizing the KV Cache can markedly reduce memory usage. Existing quantization methods have demonstrated impressive model performance even at very low bit-widths. However, these methods either rely on static one-size-fits-all precision allocation schemes (Liu et al. 2024c, a), lacking flexibility and performance-aware adaptation capabilities, or incur high computational costs of dynamic quantization while failing to adaptively prioritize critical KVs in long-context tasks (Dong et al. 2024; Duanmu et al. 2024). Therefore, they are forced to make suboptimal trade-offs among memory usage, model accuracy, and computational throughput.
To address these problems, this paper proposes KVmix, a novel mixed-precision quantization method for KV Cache. Compared to existing mixed quantization methods (Dong et al. 2024; Li et al. 2025), KVmix analyzes the importance differences of different model layers at a very low cost, thereby allowing for flexible modification of the quantization configuration based on the model’s performance requirements. This flexibility enables KVmix to maximize the compression rate of the KV Cache and the throughput of the model while maintaining controllable precision. Our specific contributions are as follows:
@@
[…]
Related Work and Motivation
Related Work
@@
To mitigate KV Cache memory challenges, researchers have developed many optimization methods, primarily centered on compression techniques and dynamic memory management (Kwon et al. 2023; Lee et al. 2024). We mainly discuss the KV Cache compression techniques related to this work. Existing compression approaches encompass quantization, sparsification (Zhang et al. 2023; Li et al. 2024), and KV Cache sharing (Sun et al. 2024; Wu and Tu 2024). Our method is orthogonal to existing weight quantization (Frantar et al. 2022; Lin et al. 2024) and sparsification methods, and it can also be used as a guideline for the importance of different layers during KV sparsification to achieve more accurate KV eviction.
@@
Extensive research has focused on reducing the memory overhead of KV Cache through quantization. For example, KIVI (Liu et al. 2024c) introduced a 2-bit asymmetric quantization technique, employing per-channel quantization for Keys and per-token quantization for Values. KVQuant (Hooper et al. 2024) proposed a non-uniform quantization strategy, integrating pre-RoPE per-channel Key quantization with per-token Value quantization. It employs offline calibration to manage outliers, achieving robust performance in long-context inference scenarios. QAQ (Dong et al. 2024) developed a dynamic mixed-precision quantization method that calculates quantization bits for Keys and Values online and optimizes the trade-off between accuracy and compression ratio by predicting attention scores. Atom (Zhao et al. 2024) investigated a mixed-precision scheme involving 4-bit and 8-bit activations, dynamically quantizing activations to adapt to input distributions. QJL (Zandieh et al. 2025) introduced a 1-bit quantization technique for the Key, leveraging the Johnson-Lindenstrauss transform followed by sign-bit quantization. KVTuner (Li et al. 2025) frames the mixed quantization of KV caches as a search optimization problem, aiming to find the optimal KV quantization configuration within a vast search space. Numerous other studies have also made significant contributions to KV Cache quantization (Yue et al. 2024; Yang et al. 2024b; Liu et al. 2024a). These efforts generally aim to quantize as many KVs as possible to the lowest feasible bit-width. While methods such as KVQuant, QAQ, and KVTuner have introduced mixed-precision quantization for KV, they often entail significant computational or search overhead. In contrast, this work introduces a lightweight and flexible framework that allows users to dynamically balance quantization bit-width and model accuracy based on specific deployment requirements.
Moreover, several works have identified the attention sink phenomenon, where attention scores excessively favor initial or recent tokens, and newly generated tokens emphasize recent contexts. StreamingLLM (Xiao et al. 2023) leverages this characteristic to achieve efficient infinite-length streaming inference by retaining attention sinks (e.g., initial tokens) alongside recent tokens. PyramidInfer (Yang et al. 2024a) applies layer-wise KV cache compression, selectively keeping key contexts based on attention patterns. Inspired by these works, we propose a dynamic pivotal context selection strategy. Unlike prior approaches, ours determines pivotal context size based on layer-specific KV importance analysis and dynamically updates full-precision KV pairs during decoding, better balancing memory efficiency and generation quality.
Motivation
Current KV Cache quantization methods rely on fixed quantization strategies, neglecting the varying contributions of KV across layers to the final output. To substantiate that quantizing Keys or Values from different layers impacts the model differently, we selectively applied 2-bit quantization to the Keys or Values of distinct layers and evaluated the effects on the model’s accuracy, with results presented in Fig. 1. The findings demonstrate that quantizing Keys or Values from different layers has varying impacts on the model’s generation quality. However, efficiently analyzing the contribution disparities across layers of the model to allocate different quantization bits to Keys or Values remains a critical challenge that needs to be addressed.
Figure 1: Accuracy of the Llama 2-7B model (Touvron et al. 2023b) on the GSM8K (Cobbe et al. 2021) and TruthfulQA (Lin et al. 2022) datasets using lm_eval (Gao et al. 2024) (FP16 represents no quantization; 0-3 indicates 2-bit quantization applied individually to the Key or Value of layers 0 through 3, respectively. And so on).
In each layer of the KV Cache, the computation process works as follows: at time step $t$, the $i$-th layer receives hidden states $H_{i-1,t}$ from the preceding layer’s output. These hidden states are used to compute the current token’s K and V as: $K_{i,t}=W_{k_{i}}\cdot H_{i-1,t}$ and $V_{i,t}=W_{v_{i}}\cdot H_{i-1,t}$, where $W_{k_{i}}$ and $W_{v_{i}}$ are the projection weights for K and V at the $i$-th layer. After computation, the computed $K_{i,t}$ and $V_{i,t}$ are concatenated with the previously stored KV, yielding the complete K and V sequences up to time step $t$: $K_{i,1:t}=[K_{i,1},K_{i,2},\dots,K_{i,t}]$ and $V_{i,1:t}=[V_{i,1},V_{i,2},\dots,V_{i,t}]$. This computation process indicates that the $W_{k_{i}}$ and $W_{v_{i}}$ determine how Key and Value are extracted from the hidden states, directly affecting the quality of the KV pairs generated by the attention mechanism and the layer’s contribution to the model’s output. Fig. 2 provides heatmaps of the $W_{k}$ and $W_{v}$ for the Llama 2-7B model, with more comprehensive heatmaps for additional models included in Technical Appendix A. The heatmaps highlight two key insights: ① Significant variations in KV weight values across different layers.Distinct distribution patterns of KV weights within the same layer. For KVs, their values dynamically adapt to changes in the input; however, $W_{k}$ and $W_{v}$ are learned during the model’s training phase and remain static during inference. As a result, these weights can be leveraged to assess the importance differences of Keys and Values across layers.
@@
Figure 2: Projection matrix weights of K and V across different layers for the Llama 2-7B model. "Norm" represents the L2 norm of weight matrix for each layer, while "Range" indicates the range of values within each layer’s weight matrix.
Methodology
KV Importance Analysis
In the attention mechanism, KV is computed by applying the KV projection weight matrices to the hidden state of the previous layer for linear transformation, then combining the result with the Query vector to calculate the attention output. Therefore, the magnitude of the KV projection weights alone is insufficient to measure the importance of KV across layers, necessitating a more precise evaluation metric. This metric should quantify the sensitivity of K and V at each layer to the model’s loss function $L$. Based on the Motivation and by using the chain rule, we can obtain:
$K_{i,t}=W_{k_{i}}\cdot H_{i-1,t}$ (1)
$\longrightarrow\nabla_{W_{k_{i}}}L=\frac{\partial L}{\partial K_{i}}\frac{\partial K_{i}}{\partial W_{k_{i}}}=\frac{\partial L}{\partial K_{i}}H_{i-1}^{T}$ (2)
[…]
$V_{i,t}=W_{v_{i}}\cdot H_{i-1,t}$ (4)
$\longrightarrow\nabla_{W_{v_{i}}}L=\frac{\partial L}{\partial V_{i}}\frac{\partial V_{i}}{\partial W_{v_{i}}}=\frac{\partial L}{\partial V_{i}}H_{i-1}^{T}$ (5)
[…]
Here, $\nabla_{W_{k_{i}}}L$ and $\nabla_{W_{v_{i}}}L$ denote the gradients of $L$ with respect to $W_{k_{i}}$ and $W_{v_{i}}$, respectively. $\frac{\partial L}{\partial K_{i}}$ and $\frac{\partial L}{\partial V_{i}}$ represent the partial derivatives of $L$ with respect to the Key and Value, respectively. To quantify the perturbation, assume that the quantization operation introduces small perturbations to the Key and Value matrices, i.e., $K_{i}^{q}=K_{i}+\Delta K$ and $V_{i}^{q}=V_{i}+\Delta V$, where $\Delta K$ and $\Delta V$ are the quantization errors. The change in $L$ due to quantization is $\Delta L=L(K^{q},V^{q})-L(K,V)$, which is the difference between the original loss and the quantized loss. To approximate $\Delta L$, perform a first-order Taylor expansion around Key:
$L(K^{q},V^{q})\approx L(K,V)+\frac{\partial L}{\partial K}\cdot\Delta K+\frac{\partial L}{\partial V}\cdot\Delta V$ (7)
Thus, the loss change due to quantization is:
$\Delta L\approx\frac{\partial L}{\partial K}\cdot\Delta K+\frac{\partial L}{\partial V}\cdot\Delta V.$ (8)
$\frac{\partial L}{\partial K}$ and $\frac{\partial L}{\partial V}$ are the gradients $\frac{\partial L}{\partial K_{i}}$ and $\frac{\partial L}{\partial V_{i}}$ computed earlier, so:
$\Delta L\approx\langle\frac{\partial L}{\partial K_{i}},\Delta K\rangle+\langle\frac{\partial L}{\partial V_{i}},\Delta V\rangle$ (9)
where $\langle\cdot,\cdot\rangle$ denotes the inner product. For a fixed $\Delta K$, a larger $\left\|\frac{\partial L}{\partial K_{i}}\right\|_{2}$ (take L2 norm) amplifies $\Delta L$, indicating greater sensitivity. The weight gradient norm thus reflects Key’s impact on $L$, as $\|\nabla_{W_{k_{i}}}L\|_{2}$ proxies $\left\|\frac{\partial L}{\partial K_{i}}\right\|_{2}$ (modulo input scaling). The Value is the same.
@@
Based on the above analysis, we propose the KVmix profiler, a gradient-based method that quantifies the contribution of each layer’s Key and Value to the model’s output, enabling a judicious mixed-precision quantization strategy. Specifically, we compute the $L2$ norm of the gradients of the model’s loss function $L$ with respect to the Key and Value projection weight matrices for each model layer ($\|\nabla_{W_{k_{i}}}L\|_{2}$ and $\|\nabla_{W_{v_{i}}}L\|_{2}$), and evaluate the importance of the Key and Value components based on these $L2$ norm values. KVmix profiler captures the dynamic sensitivity of these parameters during the model inference process, and provides a layer-specific importance metric to support efficient mixed-precision quantization in subsequent inference stages.
@@
The implementation of KVmix profiler consists of the following three key steps: ① Data preparation and forward propagation. A full-precision model is loaded, and multiple prompts are randomly sampled from a target dataset to serve as input data. These prompts are tokenized into input tensors using the tokenizer. Leveraging the autoregressive property of LLMs, each input tensor is shifted left by one position to be used as the corresponding label tensor for computing the model’s loss function. Subsequently, the loss value for each input is determined through forward propagation. ② Gradient calculation and importance assessment. For the $i$-layer of the model, the gradients of the losses with respect to the projection weights of the Key ($W_{k_{i}}$) and Value ($W_{v_{i}}$) are computed independently. This process begins with backpropagation to calculate the gradients, i.e., $\nabla_{W_{k_{i}}}L$ and $\nabla_{W_{v_{i}}}L$. The magnitude of these gradients is then evaluated using the $L2$ norm, i.e., $\|\nabla_{W_{k_{i}}}L\|_{2}$ and $\|\nabla_{W_{v_{i}}}L\|_{2}$. The importance scores of the Keys and Values for each layer can be expressed as:
$s_{k_{i}}=\|\nabla_{W_{k_{i}}}L\|_{2},s_{v_{i}}=\|\nabla_{W_{v_{i}}}L\|_{2}$ (10)
@@
A larger $s_{k_{i}}$ or $s_{v_{i}}$ signifies a greater impact of that $i$-th layer’s Key or Value on model’s output. To enhance assessment reliability, the gradient norms can be averaged across multiple prompts ($p$), yielding an average importance score for each layer’s Key and Value ($P$ is the number of prompts):
@@
We classify the importance of the Key and Value components across all model layers using the importance scores. The top 20% of the layers of $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ are quantized with high-bit representations (e.g., 3-bit or 4-bit), while the remaining 80% of layers adopt more aggressive low-bit quantization (e.g., 2-bit). This 20%-80% split is not fixed and can be dynamically adjusted according to the requirements to balance the trade-off between model accuracy and memory usage. Increasing the proportion of low-bit quantization layers can further reduce the memory usage of the KV Cache, but may sacrifice some accuracy. ③Model Configuration and Inference. The KV quantization configuration results derived from the above steps are incorporated into the model configuration, enabling the quantized model to be used for inference. The detailed workflow is depicted in Fig. 3, with the algorithmic procedure outlined in Algorithm 1 of Appendix B. The profiling is performed offline and therefore does not affect inference efficiency. Moreover, the profiling is performed once, allowing the model to reuse the results for subsequent inference tasks.
@@
Asymmetric Low-Bit Quantization
Asymmetric Quantization Strategy
We use per-channel and per-token grouping quantization methods for Key and Value, respectively. The KV Cache has the shape $[B,nh,T,D]$, where $B$ is the batch size, $nh$ is the number of attention heads, $T$ is the token sequence length, and $D$ is the head dimension. When the Key is quantized per channel ($D$), the tensor is reshaped to $B\times nh\times D,T]$, with each group comprising all tokens of a single channel. This approach is inspired by the distributional properties of the Key Cache that exhibit significant outliers in the channel dimension, i.e., certain channels exhibit significantly large magnitude values. Per-channel quantization isolates errors within each channel and prevents outliers from affecting other channels. When Value is quantized per token, the shape of the tensor is preserved, and each group contains all channels of a single token. Unlike the Key Cache, the Value Cache has no pronounced outliers, but plays a critical role in computing the attention output. Per-token quantization confines errors to individual tokens, preserving the integrity of other important tokens. This asymmetric quantization strategy effectively reduces errors introduced during KV Cache quantization.
Group-Wise Low-Bit Quantization
We use group-wise low-bit quantization to minimize KV Cache memory usage. The process includes: ① Calculation of scaling factor $s$. For each group (per-channel for Key, per-token for Value), compute $s=\frac{max\_val-min\_val}{q_{max}}$ using group min/max values, where $q_{max}$ denotes the maximum quantized value. ② Element quantization. Quantize elements with $q=round\left(\frac{x-min\_val}{s}\right)$, where $x$ represents the original element value, and $q$ is the quantized value. ③ Clipping. Limit $q$ to $\max(0,\min(q,{q_{max}}))$. ④ Storage and dequantization. The quantized values are stored using bit operations within a 32-bit integer ($int32$). For 4, 2, and 1 bit, the number of elements per $int32$ is: feat_per_int=32/bitfeat\_per\_int=32/bit. Dequantization is performed by $x=q\cdot s+min\_val$. For 3-bit quantization, we introduce a new packing strategy to maximize memory efficiency. We organize the quantized elements into blocks of 11, each stored in a 32-bit integer, with the first 10 elements quantized to 3 bits and the 11th element to 2 bits. The clipping range is adjusted based on the element index:
@@
where $i$ is the element index within a block. This strategy increases packing density by 10% over uniform 3-bit quantization that can only hold 10 elements per $int32$.
Dynamic Pivotal Context Selection